OpenStudy (s):

find dy/ dx and simplify : y=lnx/e^(2x) ?

OpenStudy (anonymous):

dy/dx= (1_2xlnx)/xe^(2x)

OpenStudy (anonymous):

or you can write it : dy/dx=(1/xe^(2x)) _2y

OpenStudy (s):

i am confused , what exactly I have to do in this problem..

OpenStudy (lgbasallote):

use quotient rule:\[\large \frac{\frac{e^{2x}}{x} + 2\ln x (e^{2x})}{e^{4x}}\]

OpenStudy (lgbasallote):

that's supposed to be minus not plus

OpenStudy (s):

ok, thanks! I just didn't knew which rule to apply

OpenStudy (lgbasallote):

QUOTIENT RULE \[\frac{ldh - hdl}{l^2}\] where l = low; h = high; d = derivative